A $400 Xbox 360 is marked down 10%. After two weeks the Xbox did not sell so it was marked down another 10%. How much is the new cost?
a. $324 b. $340 c. $385 d. $390

Answer by checkley77(9413)

400-.10*400=400-40=360-.10*360=360-36=324 ANS. A)